# “Add more config detail”: A Taxonomy of Installation Instruction Changes

### v1

@@ 5 Implications and Discussion
The ultimate goal of our study is to understand the practices of updating README files in the parts related to the installation, enabling better documentation activities to improve the quality of software documentation. In this section, we discuss the findings of this study and their implications for different stakeholders, including researchers and practitioners.
@@
For documentation maintainers, instruction completeness is important. The categories that emerged from our qualitative analysis of pre-installation instructions, installation-related instructions, and post-installation instructions collectively account for a complete process, enabling self-contained documentation for users to follow. In our synthesised codes, the introduction of a new instruction without a pre-existing alternative implies a deficiency in the previous documentation. For example, we found that the programming language and operating system to be frequently newly added to the document. This type of update addresses previously absent details. On the other hand, practitioners should also take into account the completeness of help information. Well-managed help information can improve the comprehension of the documentation and reduce the search time for users in case of errors [46]. Based on our analysed data, it is evident that the frequency of help information is notably lower compared to the combined occurrence of instructions within the three major categories. This observation implies a tendency among practitioners to overlook the inclusion of comprehensive help information.
@@
For researchers, we suggest several directions for investigation to improve the completeness of README files. First, this work primarily focuses on the update operation without looking at the state of README file before and after modification. Therefore, identifying whether the updates are complete in reflecting the system change and fixing the previous document’s lack of information would help better characterise a “good” or “bad” update. Second, an investigation of the temporal aspect of the absence of various types of instructions within the README file could better quantify the severity of this issue. Furthermore, researchers can also assist in providing essential help information. Automatically supplying help information is a difficult task. However, this can be done by algorithms to detect insufficient information for each instruction. Moreover, previous efforts [30] augmented API documentation with information from Stack Overflow. Similar ideas could be adopted by automatically extracting the most requested help information from various online channels based on certain project characteristics. Lastly, the development of tools to detect completeness issues could help document maintainers identify and address these concerns quickly.
Figure 9: Proposed README Template
GitHub could also help mitigating this issue by providing a default template for installation-related instructions when users create the new README files. A well-defined template outlines all essential aspects that should be covered in the README files.
In Figure 9, we present a suggested template based on our analysis. It covers basic instructions including prerequisites, installation steps, usage, external documents, version history, and help and support sections. In prerequisites section, we incorporate programming language version and dependencies for the project. URLs for the official websites of the dependencies are also provided. The installation steps include both source code level installation and published package level installation. Usage section covers basic commands for running the project, as well as some advanced option to display the capability of the software. External documents provides links to external comprehensive documents. Version history section provides the version information of this project, while help and support section provides most common FAQs and channels for users to interact with. In cases where specific functionalities are yet to be implemented, these placeholders can be seen as a form of self-admitted document debt, serving as reminders for document maintainers to complete the necessary information. It should be noted that we have not yet validated this template with maintainers or users.
@@
As indicated in the previous section, errors occur across different categories. Less severe errors encompass grammar issues, typographical errors within the text, and Markdown syntax errors. More severe errors are related to instructions, which require certain level of expertise to identify. Although not explicitly coded, we identified a few updates fixing errors within their instructions. In combination, these two types of errors constitute a noticeable percentage of update behaviours. This type of documentation effort can be avoided by checking the content more carefully.
Concerning considerations of up-to-dateness, delays in updating the README file can result in a temporary misalignment and potentially introduce correctness issues in the document. Codes related to updating various instructions reflect the changes in the software system. As also discovered in [47], code elements are indeed updated frequently in our analysed data. Specifically, the frequent pattern found for code update code demo for using this project corresponds to modifying code elements to reflect naming paradigm shifts within the source code. Additionally, change factors regarding meta data (project names, file path names, and technology used) of the project, modification of dependencies for the project, instructions with specific operating systems, all influence the up-to-dateness of the README file and require close monitor by document maintainers.
For researchers, an investigation of the events that trigger instructional changes in the README files could shed light on the severity of the up-to-dateness issue and reveal the relationship between coding activities and documentation activities. Second, the combination of instructional update results and trigger events can provide valuable insights for the development of automatic document tools to perform updates to mitigate up-to-dateness concerns.
@@
As indicated in Table II, document presentation undergoes frequent updates, ranking as the second most frequently updated category in our analysis. This observation underscores the considerable effort practitioners invest in updating this particular aspect of the documentation. However, most of the updates are repetitive and could be avoided by adopting better documentation schemes. Within this category, the code change section header name has the highest frequency of 88 occurrences, followed by codes that adjust the hierarchies of the sections and move the contents from one section to another. Therefore, practitioners can adopt more standardised section headers and examine whether there are mismatches between contents and headers before committing the change. For researchers, natural language processing techniques can be applied to examine content-header similarity or summarising the section contents with more appropriate headers [48]. In addition, techniques can be applied to examine and adjust the hierarchical structure for the README sections. Furthermore, this observation also motivates us to propose the README installation-related templates in Figure 9 to mitigate practitioners’ efforts in document structure modifications.
@@
Text rephrasing is also frequently observed within README updates. Two codes dedicated to adjusting the wording within the text or making sentences succinct account for 53 occurrences in total. For document maintainers, we suggest that they consider the audience of README files before writing the document. Specifically, as README files target newcomers and serve as an introductory document, overly complicated sentences and terminologies should be avoided. For researchers, previous research [49, 32] attempted to identify documentation smells or simplify complicated sentences within the documentation. With the advent of ChatGPT [50], text rephrasing becomes easier and attention is drawn to how different groups of users perceive useful information.
@@
README files serve as an introductory resource for navigating the OSS projects and should not be regarded as exhaustive documentation. In our analysed data, we observed instances where the concept of project running instructions involved presenting a comprehensive list of all running options with varying parameter combinations. This approach could lead to information overload, making it challenging for documentation users to locate their desired content. In addition, we also identified a pattern of content removal in combination with external document addition, when content within the README files becomes excessively long. For example, there are a few cases where instructions are removed, and a link to INSTALL.md or CONTRIBUTE.md is supplied in the README files.
@@
For practitioners, the above-mentioned external content organisation approach can be adopted to avoid information overload. Furthermore, we identified in code add code instruction for running this project that add a comprehensive list of project running instructions, making it difficult for users to identify the desired content. In this scenario, these instructions can be documented in a RUNNING.md with README providing a reference to it, or covered by command line interface with parameter --help.
@@
6 Threats to Validity
@@
First, we only collect repositories that use Python or Java as their primary programming language, which could affect the generalisability of our findings to other software projects. However, these two programming languages are widely-used and representative in building diverse varieties of software. The most common factors for these programming languages are shared across other languages, hence mitigating this threat.
Second, as indicated earlier, our semi-automatic labelling heuristic is not completely accurate in distinguishing “relevant” and “irrelevant” commits. In the sampled dataset, we discarded the ones that are misclassified during our manual analysis, and fill each sample bucket with additional randomly sampled commits to ensure the same amount of data is analysed.
Third, the 400 analysed repositories may not be representative to the overall population, potentially introducing sample bias. To address this threat, we ensured that sampled dataset size exceed 95% confidence rate of the overall dataset. Furthermore, our qualitative analysis may inherently carry subjectivity. To mitigate this concern, a rigorous process was implemented, involving all authors in reviewing the codes. Weekly meetings were held to facilitate discussions and address any discrepancies for codes, concepts and categories.
@@
Last, our analysis is limited in README update contents, without examining external referenced links. The results may also only apply to GitHub and are not easily generalisable to beyond GitHub projects to other platforms and closed-source projects.
7 Conclusion
@@
README files serve a critical role of introductory overviews for software projects, among which the installation-related sections are included. Considering various issues these documents may encounter, as well as the benefits of investing efforts to maintain high-quality documentation, knowledge of the practices of the README file efforts becomes essential. To achieve this, we conducted a large-scale qualitative analysis on GitHub README file commit updates specifically related to installation sections, aiming to gain a comprehensive insight into documentation activities.
@@
In this paper, we collected GitHub repositories with several stringent filters to obtain a dataset consisting of 12,908 repositories. From this dataset, we randomly sampled 400 repositories, resulting in 25,209 commits that modified a README file. We applied semi-automatic labelling heuristics and eventually obtained 8,673 commits that modified sections related to installation topics. Among these, 1,168 instances were cluster sampled for thematic analysis, resulting in a comprehensive taxonomy of 189 update behaviours.
@@
Our research reveals six categories of updates, namely pre-installation instructions, installation instructions, post-installation instructions, help information updates, document presentation, and external resource management. We further provide detailed insights into modification behaviours and offer examples of these updates.
@@
Based on our findings, we provided suggestions for practitioners and potential directions for researchers from perspectives of completeness, correctness and up-to-dateness, and information presentation considerations. We further provided a README file template tailored to cover the installation-related sections for document maintainers to get start with. To sum up in an abstract way, documentation maintainers should take care of completeness and correctness for instructions in a succinct manner, while incorporating help information and external resources to make README files more accessible and easier to comprehend. Document presentation is also a critical aspect to consider. Meanwhile, future research directions include identifying triggers for README updates, understanding user information needs in documentation, and developing automated documentation tools. Knowledge in triggering events could reveal the relationship between document and code activities, user information needs help understand the appropriate instructions to include, while all these directions contribute to the eventual automated documentation tools.
8 Replication Package
The replication package is available at https://github.com/Haoyu-Gao/README_Empirical_Study.

### v3

@@ VII Implications and Discussion
The ultimate goal of our study is to understand the practices of updating README files in the parts related to the installation, enabling better documentation activities to improve the quality of software documentation and its process. In this section, we discuss the findings of this study and their implications for different stakeholders, including researchers and practitioners.
@@
For documentation maintainers, instruction completeness is important. The categories that emerged from our qualitative analysis of pre-installation instructions, installation-related instructions, and post-installation instructions collectively account for a complete process, enabling self-contained documentation for users to follow. In our synthesised codes, the introduction of a new instruction without a pre-existing alternative implies a deficiency in the previous documentation. For example, we found that the programming language and operating system are frequently newly added to the document. This type of update addresses previously absent details. On the other hand, practitioners should also take into account the completeness of help information. Well-managed help information can improve the comprehension of the documentation and reduce the search time for users in case of errors [73]. Based on our analysed data, it is evident that the frequency of help information is notably lower compared to the combined occurrence of instructions within the three major categories. This observation implies a tendency among practitioners to overlook the inclusion of comprehensive help information. Therefore, practitioners could closely monitor the issues and pull requests to incorporate frequently-asked or confused questions into the README documents. For documentation updates, they could use our proposed template as a checklist before performing the updates.
@@
For researchers, we suggest several directions for investigation to improve the completeness of the README files. First, this work primarily focuses on the update operation with less attention on the state of README file before and after modification. Therefore, identifying whether the updates are complete in reflecting the system change and fixing the previous document’s lack of information would help better characterise a “good” or “bad” update. Second, an investigation of the temporal aspect of the absence of various types of instructions within the README file could better quantify the severity of this issue. Furthermore, researchers can also assist in providing essential help information. Automatically supplying help information is a difficult task. However, this can be done by algorithms to detect insufficient information for each instruction. Moreover, previous efforts [33] augmented API documentation with information from Stack Overflow. Similar ideas could be adopted by automatically extracting the most requested help information from various online channels based on certain project characteristics. Lastly, the development of tools to detect completeness issues could help document maintainers identify and address these concerns quickly.
GitHub could also help mitigate this issue by providing the template proposed in this study for installation-related instructions when users create or update the README files. According to the result discussed in the previous section, by simply applying the template to the original document, documentation readers could experience increased documentation quality. In fact, the survey participants left comments on our proposed template: “Good to see that you provide Frequently Asked Questions (FAQ)”, “Covers all the important elements”. Therefore, this template could serve as a good starting point for reminding practitioners of the essential aspects to cover for their README files before performing updates.
@@
As indicated in the previous section, errors occur across different categories. Less severe errors encompass grammar issues, typographical errors within the text, and Markdown syntax errors. More severe errors are related to instructions, which require a certain level of expertise to identify. Although not explicitly coded, we identified a few updates fixing errors within their instructions. In combination, these two types of errors constitute a noticeable percentage of update behaviours. This type of documentation effort can be avoided by checking the content more carefully.
Concerning considerations of up-to-dateness, delays in updating the README file can result in a temporary misalignment and potentially introduce correctness issues in the document. Codes related to updating various instructions reflect the changes in the software system. As also discovered in [74], code elements are indeed updated frequently in our analysed data. Specifically, the frequent pattern found for code update code demo for using this project corresponds to modifying code elements to reflect naming paradigm shifts within the source code. Additionally, change factors regarding metadata (project names, file path names, and technology used) of the project, modification of dependencies for the project, and instructions with specific operating systems, all influence the up-to-dateness of the README file and require close monitoring by document maintainers.
@@
As indicated in Table II, document presentation undergoes frequent updates, ranking as the second most frequently updated category in our analysis. This observation underscores the considerable effort practitioners invest in updating this particular aspect of the documentation. However, most of the updates are repetitive and could be avoided by adopting better documentation schemes. Within this category, the code change section header name has the highest frequency of 88 occurrences, followed by codes that adjust the hierarchies of the sections and move the contents from one section to another. Therefore, practitioners can adopt more standardised section headers and examine whether there are mismatches between contents and headers before committing the change. For researchers, natural language processing techniques can be applied to examine content-header similarity or summarise the section contents with more appropriate headers [75]. In addition, techniques can be applied to examine and adjust the hierarchical structure for the README sections. Furthermore, this observation also motivates us to propose the README installation-related templates in Figure 10 to mitigate practitioners’ efforts in document structure modifications.
@@
Text rephrasing is also frequently observed within README updates. Two codes dedicated to adjusting the wording within the text or making sentences succinct account for 53 occurrences in total. For document maintainers, we suggest that they consider the audience of README files before writing the document. Specifically, as README files target newcomers and serve as an introductory document, overly complicated sentences and terminologies should be avoided. For researchers, previous research [76, 35] attempted to identify documentation smells or simplify complicated sentences within the documentation. With the advent of ChatGPT [77], text rephrasing becomes easier, and attention is drawn to how different groups of users perceive useful information.
@@
README files serve as an introductory resource for navigating the OSS projects and should not be regarded as exhaustive documentation. In our analysed data, we observed instances where the concept of project running instructions involved presenting a comprehensive list of all running options with varying parameter combinations. This approach could lead to information overload, making it challenging for documentation users to locate their desired content. In addition, we also identified a pattern of content removal in combination with external document addition, when content within the README files becomes excessively long. For example, there are a few cases where instructions are removed and a link to INSTALL.md or CONTRIBUTE.md is supplied in the README files.
@@
For practitioners, the above-mentioned external content organisation approach can be adopted to avoid information overload. Furthermore, we identified in code add code instruction for running this project that add a comprehensive list of project running instructions, making it difficult for users to identify the desired content. In this scenario, these instructions can be documented in a RUNNING.md with README providing a reference to it, or covered by a command line interface with parameter --help.
@@
Automatic documentation updates. By monitoring common open-source channels such as commit messages, pull requests, issues, and previous versions of the documentation, we identified various triggers for documentation updates. An automatic tool for suggesting documentation updates by taking this information into consideration could help developers quickly identify the need for documentation updates during the software development process. Future research could also investigate more from detailed source code level to identify additional triggers for documentation updates.
Additionally, the results in the template validation section indicate that documentation users perceive improvements in four key dimensions – structure, understandability, conciseness, and readability – after applying our template to the original README documents. In short, the application of this template directly mitigates the above-mentioned issues related to documentation structure, targeted audience, and information overload. In particular, one participant commented on the templates “I like the order of chapters and it is easy to find everything. Just enough information to get started with technology.”. In addition, our pull request study further shows that applying this template is also perceived as useful by documentation maintainers.
Therefore, to enhance the downstream users’ experience, documentation maintainers could manually adopt our template, or use an automatic approach of delegating this task to LLMs. The automatic approach could further alleviate the massive efforts from practitioners to adjust documentation presentation. However, this process should incorporate a human-in-the-loop approach, involving manual review and adjustments to account for specific details, as a general template may not be suitable for all scenarios.
For researchers, our approach did not extensively adopt prompt engineering or make adjustments to the generated documentation. Refinements in these aspects might further improve the quality of README documents. The usability of the augmented documents could be verified through live installation sessions with documentation users.
Furthermore, integrating documentation update suggestion tools with our proposed template has the potential to automate documentation updates, addressing issues related to up-to-dateness, correctness, and presentation simultaneously. This pipeline could be seamlessly integrated into the GitHub pull request process, where documentation suggestions are generated based on changes in relevant channels and the status of the previous version of the README. Future research could focus on refining these tools or enhancing the template to realise this vision.
VIII Threats to Validity
@@
First, we only collect repositories that use Python or Java as their primary programming language, which could affect the generalisability of our findings to other software projects. However, we did not observe any significant differences in the taxonomy between Python and Java projects in our analyzed data. The taxonomy derived in this study includes modifications in the installation process, help information, documentation presentation, and resource management. These factors can serve as a starting point for researchers and practitioners when addressing projects with different programming languages.
Second, as indicated earlier, our semi-automatic labelling heuristic is not completely accurate in distinguishing “relevant” and “irrelevant” commits. In the sampled dataset, we discarded the ones that were misclassified during our manual analysis, and filled each sample bucket with additional randomly sampled commits to ensure the same amount of data is analysed.
Third, the 400 analysed repositories may not be representative of the overall population, potentially introducing sample bias. To address this threat, we ensured that the sampled dataset size exceeded a 95% confidence rate with a 5% margin of error for the overall dataset. Furthermore, our qualitative analysis may inherently carry subjectivity. To mitigate this concern, a rigorous process was implemented, involving all authors in reviewing the codes. Weekly meetings were held to facilitate discussions and address any discrepancies in codes, concepts and categories.
@@
Fifth, our analysis is limited in README update contents, without examining external referenced links. The results may also only apply to GitHub and are not easily generalisable beyond GitHub projects to other platforms and closed-source projects.
Sixth, our procedure for identifying triggers for documentation updates may miss certain triggers, such as changes embedded in commits outside of pull requests. Additional reasons for updates might emerge from a deeper analysis of the codebase. However, the procedure was designed to monitor the primary channels within open-source repositories. This approach provides a systematic way to identify triggers without requiring exhaustive manual searches. Moreover, this design offers a foundation for automating the linkage between triggers and documentation updates, potentially enabling the identification of more triggers in the future.
Seventh, our template validation was conducted on 20 README documents with 30 participants, which might not represent the entire range of README documentation genres or the broader participant population. To mitigate this limitation, we randomly sampled the evaluated documents and recruited participants who work in IT with an additional “quality gate”. However, the participants were not actual README users; they were only required to evaluate the perceived documentation quality without installing the software according to the guidelines provided in the documents. Additionally, no formal training of participants except the definition is provided before the survey, leading to a relatively lower agreement with some level of subjectivity. Therefore, we only claim an improvement in the documentation reading experience rather than in documentation accuracy. Future work can further investigate this perspective by involving actual README users with live installation sessions.
Lastly, the pull request validation study only covers 30 repositories and may not generalise to all repositories on GitHub. Moreover, this investigation does not provide long-term data about whether the merged changes will remain in the repository. Furthermore, some repositories had not responded to our pull requests at the time of paper submission, preventing us from including those results in the current version of the manuscript.
IX Conclusion
@@
README files serve a critical role of introductory overviews for software projects, among which the installation-related sections are included. Considering various issues these documents may encounter brought by the rapidly evolving software repositories and ecosystem, as well as the benefits of investing efforts to maintain high-quality documentation, knowledge of the practices of the README file efforts becomes essential. To achieve this, we conducted a large-scale qualitative analysis on GitHub README file commit updates and their associated triggers specifically related to installation sections, aiming to gain comprehensive insight into documentation activities.
@@
In this paper, we collected GitHub repositories with several stringent filters to obtain a dataset consisting of 12,908 repositories. From this dataset, we randomly sampled 400 repositories, resulting in 25,209 commits that modified a README file. We applied semi-automatic labelling heuristics and eventually obtained 8,673 commits that modified sections related to installation topics. Among these, 1,168 instances were cluster sampled for qualitative analysis, resulting in a comprehensive taxonomy of 189 update behaviours.
@@
Our research reveals six categories of updates, namely pre-installation instructions, installation instructions, post-installation instructions, help information updates, document presentation, and external resource management. We further provide detailed insights into modification behaviours and offer examples of these updates. We also investigated the triggers for documentation updates, which led to three categories including errors in the previous documentation, changes in the codebase, and need for documentation improvement.
@@
Based on our findings, we proposed a README template for documentation maintainers when updating documents. The template validation process finds that the augmented documents based on our template are perceived as higher quality by documentation readers. Moreover, our pull request study also shows that documentation maintainers find it useful. This step serves as an initial step towards automated documentation updates to facilitate better documentation activities. Furthermore, we provided suggestions for practitioners and potential directions for researchers from perspectives of completeness, correctness and up-to-dateness, and information presentation considerations. To sum up in an abstract way, documentation maintainers should take care of completeness and correctness for instructions in a succinct manner, while incorporating help information and external resources to make README files more accessible and easier to comprehend. Document presentation is also a critical aspect to consider. Meanwhile, future research directions include identifying triggers for README updates, understanding user information needs in documentation, and developing automated documentation tools. Knowledge of triggering events could reveal the relationship between document and code activities, user information needs help understand the appropriate instructions to include, while all these directions contribute to the eventual automated documentation tools.